OpenStudy (anonymous):

Will fan and medal! The equation of the line of best fit of a scatter plot is y = −5x − 2. What is the slope of the equation? −5 −2 2 5

OpenStudy (anonymous):

The slope is the number followed by the "x"

OpenStudy (r.hamza17):

As the equation can be written as, -5x-y-2=0 so using formula for slope which is slope=-b/a slope=-(-5/1)=5